OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The third step would be to: Select one: a. Graph the function. b. Take the log of each side. c. Square the expression. d. Factor the resulting quadratic equation.

OpenStudy (solomonzelman):

log_a (b) + log_a (c) = log_a (b times c)

OpenStudy (anonymous):

so the answer B would be the first step right

OpenStudy (solomonzelman):

Nope

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log_2x + \log_2(x – 7) = 3\LARGE\color{white}{ \rm │ }}\) Step 1. \(\normalsize\color{blue}{ \log_2~x(x – 7) = 3\LARGE\color{white}{ \rm │ }}\) Step 2. \(\normalsize\color{blue}{ \log_2~x^2-7x = 3\LARGE\color{white}{ \rm │ }}\) Step 3. \(\normalsize\color{blue}{ 2^3=x^2-7x\LARGE\color{white}{ \rm │ }}\) Step 4. \(\normalsize\color{blue}{ 9=x^2-7x\LARGE\color{white}{ \rm │ }}\) Step 5. \(\normalsize\color{blue}{ 0=x^2-7x-9\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

This is what I would do to solve it.

OpenStudy (anonymous):

So step 2 is Square the expression correct

OpenStudy (loser66):

it's now square, It's property of log

OpenStudy (solomonzelman):

No, I am multiplying inside the log.

OpenStudy (solomonzelman):

but then, I am using a "log-relation" (or whatever it is actually called)

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

So b is the answer then @SolomonZelman

OpenStudy (solomonzelman):

You aren't taking the log of each side, are you ?

OpenStudy (solomonzelman):

or if you do, you get "log of log" on the left.

OpenStudy (anonymous):

So what would it be @SolomonZelman

OpenStudy (solomonzelman):

It is hardly any of the choices. Not A, because there is no need graphing to find the solution. Not B, because you aren't logging both sides. C -- don't get it. D, is it. 2^3 = 8 (not 9, sorry)

OpenStudy (solomonzelman):

So you would get x^2-7x-8=0 (x-8)(x+1)=0 x=8,-1

OpenStudy (solomonzelman):

Sorry one more time... makes sense now ?